OpenStudy (anonymous):

A 15 foot ladder is leaning against a building. The base of the ladder begins to slide away from the wall at a rate of 2.5 ft/s

OpenStudy (anonymous):

b) at what rate is the angle formed by the ladder and the floor decreasing when the base of the ladder is 10 ft from the wall?

OpenStudy (anonymous):

pythag for this one \[x^26y^2=15^2\\2xx'+2yy'=0\] or \[xx'+yy'=0\]

OpenStudy (anonymous):

oops now we are in to trig

OpenStudy (anonymous):

i have no idea how to start it off...

OpenStudy (anonymous):

draw use tangent

OpenStudy (anonymous):

or sine maybe

OpenStudy (anonymous):

or cosine

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\cos(\theta)=\frac{x}{15}\]

OpenStudy (anonymous):

isnt x 10 though?

OpenStudy (anonymous):

x is a variable

OpenStudy (anonymous):

you are looking for the rate of change of \(\theta\) right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

start with \[\huge \cos(\theta)=\frac{x}{15}\] sp \[\huge -\sin(\theta)\theta'=\frac{x'}{15}\] by the chain rule

OpenStudy (anonymous):

ok so thats exaclty where im confused where did sine come from

OpenStudy (anonymous):

it is minus sine

OpenStudy (anonymous):

on account of that is the derivative of cosine right?

OpenStudy (anonymous):

it is? i haven't done derivatives with trig yet so..... highschool calculus if your wondering

OpenStudy (anonymous):

you cannot do this without trig no way, no how

OpenStudy (anonymous):

oh okay so maybe i didnt have to do this part of the question then... since i havent learned how to derive trig yet